what is the vapor pressure of water above a solution prepared by dissolving 28.5 g glycerin (C_(3)H_(8)O_(3)) in 0.125 kg water at 243 K. The vapor pressure of pure water at that temperature is 233.70 torr.


1
Expert's answer
2021-12-15T10:17:58-0500

From Raoult's law, "P_{solution}=X_{solvent}\u00d7P_{solvent}"

Where denotes vapor pressure, and X represent mole fraction.

"Mole fraction =\frac{moles of solvent}{moles of solute+moles of solvent}"

"Vapor pressure=\frac{\frac{0.125×1000}{18}}{\frac{28.5}{92}+\frac{0.125×1000}{18}}×233.7=223.72"